OpenStudy (anonymous):

Find the limit as x approaches 0 for sinx/(5√x)

OpenStudy (zehanz):

You could use l'Hôpital's Rule: because the limit has the form x => 0 sinx/5√x => 0/0, you may use that rule, so the limit, if it exists is equal to the limit of the quotient of the derivatives :\[\frac{ 1 }{ 5 } \lim_{x \rightarrow 0}\frac{ \cos x }{ \frac{ 1 }{ 2\sqrt{x} } }\]If you simplify the fraction, you will see where this is going to...

OpenStudy (anonymous):

i dont understnd where u go from that fraction....